OpenStudy (anonymous):

Solve for x. Answer No solution 0 ±2 ±3

OpenStudy (anonymous):

 @jennychan12

OpenStudy (anonymous):

@jennychan12 help please

OpenStudy (jennychan12):

3x+1 = x^2+2x+1 x^2 -x = 0 x(x-1) = 0 x = 0, 1 your only choice is 0

OpenStudy (anonymous):

I checked my work and got no solution either that or 0